Rabbits in a lab are to be kept on a strict daily diet that includes 30 grams of protein, 16 grams of fat and 24 grams of carbohydrates. the scientist has only three food mixes available with the following grams of nutrients per unit.
Protein fat carbohydrate
mix A 4, 6, 3
mix B 6, 1, 2
mix C 4, 1, 12
Find how many units of each mix are needed daily to meet each rabbits dietary need.
(Linear equations)

Answer by jim_thompson5910(35256)

Hint:

Let
x = amount of mix A (in units)
y = amount of mix B (in units)
z = amount of mix C (in units)

So if the diet "includes 30 grams of protein, 16 grams of fat and 24 grams of carbohydrates", and the mixes protein, fat and carb counts are given in the table, then this means

4x + 6y + 4z = 30 (I got this by reading down the first column of the table, then setting it equal to the protein count)

6x + 1y + 1z = 16 (same thing as above, but this time I'm in column 2)

3x + 2y + 12z = 24

-------------------------------------------------------

So after everything is all set up, you'll have these 3 equations

4x + 6y + 4z = 30
6x + 1y + 1z = 16
3x + 2y + 12z = 24